In the following table the activities with their precedence sequence and expected time​ (days) are listed for a project on which Carl​ Betterton's construction company is​ working:

                                                                                      

Activity

Immediate​ Predecessor(s)

Time​ (days)

A

minus−

5

B

A

4

C

A

6

D

B

6

E

B

4

F

C

4

G

D

6

H

​E, F

7

​Carl's team should be able to complete the project in

nothing

days.

Critical path activities for​ Carl's project are

.
Answer #1

The various paths and their expected completion time of the project are

  • A-B-D-G = 5+4+6+6 = 21 days
  • A-B-E-H = 5+4+4+7= 20 days
  • A-C-F-H = 5+6+4+7 = 22 days

So the critical path is A-C-F-H as it has the longest expected completion time of 22 days

Expected completion time of the project = expected completion time of the critical path = 22 days. So Carl's team should be able to complete the project in 22 days